forgetmylife0:

Two friends, Rakesh and Vishal started climbing a pyramid-shaped hill. Rakesh climbs 315 m and finds that the angle of depression is 72.3 degrees from his starting point. How high is he from the ground?

Dva:

Dva:

To solve m we could use the sine ratio Sin 72.3° = m/315 0.953 = m/315 m= 315 x 0.953 Would equal what?

forgetmylife0:

300.195

Extrinix:

Correct!